OpenStudy (anonymous):

How would you solve the integral sqrt(a^2-x^2) from -a to a in the simplest easiest way ? =)

OpenStudy (lgbasallote):

if i recall right....there was this law that says \[\int \limits_{-a}^a (\text{even fuction} )dx = 0\] or was it for the odd function... i'll try and check

OpenStudy (lgbasallote):

hmm lol seems it's for the odd function

OpenStudy (anonymous):

lol !! but at least you reminded me this rule as well ;)

OpenStudy (anonymous):

the answer should be pi*a^2/2 but i can't get it right =\

OpenStudy (lgbasallote):

ahh i got it

OpenStudy (lgbasallote):

\[\Large \int\limits_{-a}^a (a^2 - x^2)dx \implies 2\int\limits_{-a}^a x^2 dx\] does that help?

OpenStudy (anonymous):

you mean twice the integral from 0 to \(a\)

OpenStudy (anonymous):

oh it is the square root!!!

OpenStudy (anonymous):

\[\int_{-a}^a\sqrt{a^2-x^2}dx\] like that i bet

OpenStudy (anonymous):

do not integration at all

OpenStudy (anonymous):

\(y=\sqrt{a^2-x^2}\) represents a circle, center at the origin, radius of \(a\) geometry tells you that the area of the circle is \(\pi a^2\) since you are taking half of that circle, area is \[\frac{\pi a^2}{2}\]

OpenStudy (anonymous):

i meant of course the upper half of the circle 

OpenStudy (anonymous):

Great !!!!! thank you !!!!!!! =) Nice thinking !!!